OpenStudy (anonymous):

How do you simplify this equation (8+√5)(8-√5)?

jimthompson5910 (jim_thompson5910):

Use the idea that \[\Large (x+y)(x-y) = x^2 - y^2\]

OpenStudy (anonymous):

I don't get that..

jimthompson5910 (jim_thompson5910):

Compare \[\Large (x+y)(x-y)\] with \[\Large (8+\sqrt{5})(8-\sqrt{5})\]

jimthompson5910 (jim_thompson5910):

Can you see how x = 8 and \(y = \sqrt{5}\) ???

OpenStudy (anonymous):

Right.

jimthompson5910 (jim_thompson5910):

So \[\Large (x+y)(x-y) = x^2 - y^2\] becomes \[\Large (8+\sqrt{5})(8-\sqrt{5}) = (8)^2 - (\sqrt{5})^2\]

OpenStudy (richyw):

just expand it!\[(8+\sqrt{5})(8-\sqrt{5})\] \[=64-8\sqrt{5}+8\sqrt{5}-5\] \[=64-5=59\]

hero (hero):

Difference of Squares Formula: \(a^2 - b^2 = (a+b)(a-b)\)

hero (hero):

In this case, \(a = 8, b = \sqrt{5}\)

hero (hero):

So by definition of difference of squares: \(a^2 - b^2 = (a+b)(a-b)\) \(8^2 - \sqrt{5}^2 = (8 + \sqrt{5})(8-\sqrt{5})\) \(64 - 5 = (8 + \sqrt{5})(8-\sqrt{5})\) \(59 = (8 + \sqrt{5})(8-\sqrt{5})\)

hero (hero):

Proof of squares: \(a^2 - b^2 = (a+b)(a-b)\) \(=a(a-b)+b(a-b))\) \(=a^2 - ab + ab - b^2\) \(=a^2 - b^2\)

hero (hero):

Further proof: